Helping my grandson with 6th grade Algebra.
x+3x=x+x+10
How do I break down to get started in solving?

Answer by rfer(16322)

combine all of the "x"
x+3x=x+x+10
4x=2x+10
subtract 2x from each side
4x-2x=2x-2x+10
2x=10
divide both by 2
2x/2=10/2
x=5